OpenStudy (anonymous):

A yield sign is an equilateral triangle, with side length of 34 inches. How many square inches of material are needed to make a sign? Round your answer to the nearest tenth of a square inch. (Points : 2) 499.8 in2 578 in2 289 in2 705.6 in2

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

A = (1/4)*(s^2)*sqrt(3) A = (1/4)*(34^2)*sqrt(3) A = (1/4)*(1156)*sqrt(3) A = ???

OpenStudy (anonymous):

thts hard

jimthompson5910 (jim_thompson5910):

I'm getting 500.56268 which isn't an answer choice, but it's close to choice A so it might be due to rounding error

OpenStudy (anonymous):

289

OpenStudy (anonymous):

look wat i found but i dnt think they giv her the answer http://openstudy.com/study#/updates/500940c5e4b020a2b3be6202

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

yeah they're saying the same thing

OpenStudy (calculusfunctions):

There is more than one way to view this problem. First you know that the area of a triangle is Area = (0.5)(base)(height). In any equilateral triangle\[A =\frac{ \sqrt{3} }{ 2 }b ^{2}\]Since all three sides of a triangle are known, you could also apply Heron's formula for the area of a triangle, which is\[A =\sqrt{s(s -a)(s -b)(s -c)}\]where s is half of the perimeter of the triangle and a, b, and c are the side lengths.

OpenStudy (calculusfunctions):

Sorry, that should be\[A =\frac{ \sqrt{3} }{ 4 }b ^{2}\]